# Automatically Interpreting Millions of Features in Large Language Models

### v2

@@ 3.3.1 Detection
One simple approach for scoring interpretations is to ask a language model to identify whether a whole sequence activates a SAE latent given an interpretation. Detection requires few output tokens for each example used in scoring, meaning examples from a wider distribution of latent activation strengths can be used at approximately the same expense. By including non-activating contexts, this method measures both the precision and recall of the interpretation.
Detection is more “forgiving” than simulation insofar as the scorer does not need to localize the latent to a particular token. This means that detection evaluates whether an interpretation correctly identifies the types of contexts that a latent is active on, even if the interpretation does not correctly identify the actual token.
Both this method and fuzzing (described next) can leverage token probabilities to estimate how certain the scorer model is of their classification, and this is an effect that we believe can be to improve the scoring methods. Details on the prompt in Appendix A.4.1.
@@ 3.3.2 Fuzzing
When interpretations focus on which tokens the latent activates, but not on which contexts, they can still score high on fuzzing, but comparatively lower on detection, specially if they are common tokens. Thus, evaluating an interpretation on both detection and fuzzing can identify whether a model is classifying examples for the correct reason. Details on the prompt on Appendix A.4.2.
@@ 3.3.3 Surprisal
Surprisal scoring is based on the idea that a good interpretation should help a base language model $\mathcal{M}$ (the “scorer”) achieve lower cross-entropy loss on activating contexts than it would without a relevant interpretation. Specifically, for each context $\mathbf{x}$, activating or non-activating, we measure the information value of an interpretation $\mathbf{z}$ as $\log p_{\mathcal{M}}(\mathbf{x}|\mathbf{z})-\log p_{\mathcal{M}}(\mathbf{x}|\tilde{\mathbf{z}})$, where $\tilde{\mathbf{z}}$ is a fixed pseudo-interpretation. A good interpretation should have higher information value on activating examples than on non-activating ones.
The overall surprisal score of the interpretation is given by the AUROC of its information value when viewed as a classifier distinguishing activating from non-activating contexts. Details on the prompt and how to compute the score in Appendix A.4.3.
@@ 3.3.4 Embedding
Classifying between active and non-active contexts given a certain interpretation can also be seen as using interpretations of latents as “queries” that should be able to retrieve relevant “documents”, contexts where the latent is active, between non-relevant “documents”, non-activating contexts. This way, we take a selection of activating and non-activating contexts that embedded by an encoding transformer, and the similarity between the query and the documents is used as a classifier to distinguish between activating and non-activating contexts, and the score is given by the AUROC.
If the encoding model is small enough - we used a 400M parameter model - this technique is the fastest and opens up the possibility to evaluate a larger fraction of the activation distribution. We have seen that using a larger embedding model - 7B parameter model - didn’t significantly improve the scores, see fig A3, although we believe that this approach was under-investigated. Details on the prompt, on the embedding model and on the way to compute the score in Appendix A.4.4.
@@ 3.3.5 Intervention scoring
Unlike the above four context-based scores, intervention scoring interprets a latent’s counterfactual impact on model output. We quantify the interpretability of an intervention $I$ with interpretation $\mathbf{z}$ on a distribution of prompts $\pi$ as the average decrease in the scorer’s surprisal about the interpretation when conditioned on text generated with the intervention.
$S(I,\mathbf{z};\pi)=\mathbb{E}_{\mathbf{x}\sim\pi}\Big{[}\mathbb{E}_{\mathbf{i}\sim\mathcal{G}_{I}(\mathbf{x})}[\log p_{\mathcal{M}}(\mathbf{z}|\mathbf{i})]-\mathbb{E}_{\mathbf{g}\sim\mathcal{G}(\mathbf{x})}[\log p_{\mathcal{M}}(\mathbf{z}|\mathbf{g})]\Big{]}$ (1)
@@
$\sigma(I;\pi)=\mathbb{E}_{\mathbf{x}\sim\pi}\big{[}D_{KL}(p_{\text{subject}}(\cdot|\mathbf{x})\ ||\ p_{\text{subject},I}(\cdot|\mathbf{x}))\big{]}$ (2)
See Appendix A.7 for details on the interpretation and scoring pipeline we use in our intervention experiments.
@@ 4.1 Comparing scoring methods
When wanting to evaluate the explanation over a larger number of examples, simpler methods than simulation scoring have to be used (Templeton et al., 2024). Simulation scoring is traditionally done only in activating examples, ignoring whether the proposed explanation correctly handles non-activating examples. Simulation scoring is normally done in a small amount of examples per latent due to its high cost, and this is a significant disadvantage that prompted us to investigate more efficient scoring methods.
In Table 1 we compare the amount of tokens used to score a single latent using 100 different contexts, for both fuzzing, detection and two different simulation methods. We can see that fuzzing and detection are at least 5x cheaper than simulation when using the all-at-once (AAO) method described in Bills et al. (2023), although this method requires access to the log-probabilites of the prompt tokens, which are not accessible from providers of state-of-the-art closed-source models. A researcher wanting to generate simulation scores on the fly, without access to a local model, would have to use the token-by-token (TBT) method, which is 30x more than expensive than both fuzzing and detection. For the same set of intepretations, fuzzing and detection scores given by Llama 70b and Claude Sonnet 3.5 have similar distributions. On the other hand, Claude Sonnet 3.5 produces to higher simulation scores on average than Llama 70b, see Table A9 meaning that we should probably compare the price of fuzzing and detection on Llama 70b to the price of simulation using Claude Sonnet 3.5.
Embedding scoring is even cheaper, requiring 4000 input tokens for 100 different contexts, which at an average of \$0.13 per million tokens would only cost \$50/100k latents.
@@
Input tokens Output tokens Cost (\$/100k latents) Llama 70b Cost (\$/100k latents) Claude Sonnet 3.5
@@
We propose that latent interpretations should be evaluated using more than a single technique when possible, because as discussed in Section 3.3, each of the proposed methods have different failure modes, and a more rich scoring framework can address some of these flaws. Specifically, we find that when computing the correlation of scores computed on the same 800 explanations, we find that while there is a clear positive correlation, some scores disagree – see Appendix A.4.6 and Tables A1 and A2.
@@
We find fuzzing to be the most correlated with simulation, as fuzzing is mostly measuring how well the explanation can help the model predict which tokens have non-zero activation. Due to the sparsity of SAE activations, simulation scoring is doing a very similar test. On the contrary, we find that simulation has a significantly lower correlation with detection, embedding and surprisal scoring, as they more accurately measure whether a context activates a given latent, instead of which token is active in a given context. When comparing the simulation scores provided by Claude 3.5 Sonnet, all correlations have a slight increase, see Appendix A.4.6.
@@
We also measure the correlation between our scoring methods and scores given by humans. Humans were tasked to score how an explanation related to a given context, similar to that done in (Templeton et al., 2024). The human evaluators could choose between 4 options, “1 - Explanation not related to the text”, “2 - Explanation vaguely related to the text”, “3 - Incomplete explanation but present in the text”, “4 - Explanation describes the activation of the tokens”. We measure the average score of a latent in at least 5 contexts. In total 700 contexts and 81 latent interpretation. We find that fuzzing as the best Spearman correlation with human scores (0.69), followed by simulation (0.60), and detection (0.59). Surprisal and embedding have 0.34 and 0.32 Spearman correlation respectively.
@@
Due to their relative cheapness compared with simulation scoring, we propose that fuzzing and detection scoring can be used to estimate whether explanations correctly identify on which tokens a latent is most likely to be active, and in which types of contexts this happens. Cheaper methods like embedding can be used to quickly iterate on explanations or to broadly separate latents that have bad explanations from those with good explanations, so that they can be refined.
@@ 4.1.1 Intervention scoring
The previously discussed scoring methods are correlational in the sense that they measure whether an interpretation can be used to predict the activation values of a given latent, or distinguish between activating and non-activating examples. Intervention scoring proposes to measure how well a given interpretation can predict the effect of interventions on the corresponding latent. Here we compare correlational interpretations generated with our pipeline and scored with fuzzing, to a set of interventional interpretations scored with intervention scoring. Our hypothesis is that some latents will have low correlational scores because their behavior is better explained by their downstream effects than by the contexts where they are active.
In Figure 4, we see that there is a slight negative correlation between the fuzzing score and the intervention score, showing that on average, latents with low fuzzing score can be better explained using their downstream effects than latents with high fuzzing score. We also find that the distribution of intervention scores of interpretation on a trained SAE is significantly different from the distribution of an interpretation of a randomly initialized SAE or a real SAE with randomly assigned interpretations, supporting the validity of this scoring method.
@@ 4.2 Comparing interpretation methods
We use a set of 500+ latents as a testbed to measure the effects of design choices and hyperparameters on interpretation quality. Each latent is scored using 100 activating and 100 non-activating examples. The activating examples are chosen via stratified sampling such that there are always 10 examples from each of the 10 deciles of the activation distribution. We evaluate interpretation quality using fuzzing, detection and embedding scores, as those were both quick to compute and easy to interpret. We expect this mix of scores reflect the extent to which the proposed interpretation is valid over both activating and non-activating examples.
@@
We find that the interpretations generated by considering only the activating contexts are significantly different from those generated by selecting activating contexts from the whole distribution. Interpretations based on top examples have higher specificity, but lower sensitivity. In fact, we observe that the sensitivity depends on how the degree of activation of each example shown, and that this sensitivity decreases faster for interpretation based on top examples. Interpretations generated from top activating examples are more concise and specific interpretations but fail to capture the whole distribution, while explanations based on examples sampled from the whole distribution lead to interpretations that are too broad, see Appendix A.3. This effect would not be seen if the scoring were done on just the most activating examples, underscoring a problem with current auto-interpretability evaluations, which produce interpretations using top activating examples and evaluate them on a small subset of the activation distribution.
@@
Increasing the size of the explainer model increases the scores of the interpretations, but we don’t find the explanations generated by Claude Sonnet 3.5 to have much higher scores than those generated by Llama 3.1 70b, see A9, and both are similar to those generated by a human. Not surprisingly, we also see that even for simpler scores like fuzzing and detection, using a smaller scorer model leads to lower scores. It is possible that the explanations generated by Claude would be better than those generated by Llama 3.1 70b had we optimized the prompting techniques for that model.
@@
Showing the explainer model a larger number of examples leads to slightly higher scores (Table A7). We find it doesn’t matter much whether we use the same dataset as the training set of the SAEs (Table A3) or whether we slightly change the size of the contexts shown (Table A5). Using COT, at least when generating interpretations using Llama 70b, does not significantly increase their quality (Table A4) while significantly increasing the compute and time required to generate them. For this reason, we have not used it for our main experiments.
@@
We find that SAEs with more latents have higher scores, and scores that are significantly higher than those of neurons. Neurons are more interpretable if made sparser by only considering the top $k$ most activated neurons on a given token, but still significantly underperform SAEs in our tests (Table A10). The location of the SAE matters; residual stream SAEs have slightly better scores than ones trained on MLP outputs (Table A10). We also observe that earlier layers have lower overall scores, but that the scores across model depth remain constant after the first few layers (Fig A5).
Figure 4: Intervention scores. Here we present intervention scores (Sec 3.3.5) for SAE latents in Gemma 2 9B at layer 32. Left: SAE latents are more interpretable than random latents, especially when intervening more strongly. Our explainer also produces interpretations that are scored higher than random interpretations. Right: Many latents that would normally be uninterpreted when using context-based automatic interpretability are interpretable in terms of their effects on output.
@@ 5 Conclusion
Explaining the latents of SAEs trained on cutting-edge LLMs is a computationally demanding task, requiring scalable methods to both generate interpretations and assess their quality. We addressed issues with the conventional simulation-based scoring and introduced five new scoring techniques, each with distinct strengths and limitations. These methods allowed us to explore the “prompt design space” for generating effective interpretations and propose practical guidelines.
Additionally, although current scoring methods do not account for interpretation length, we believe shorter interpretations are generally more useful and will incorporate this in future metrics. Some scoring methods also require further refinement, particularly in selecting non-activating examples to improve evaluation. We also investigated a method to generate and scoring interpretations of latents that is based on their downstream effect and shown that low score “correlational” scores could be due to the existence of “output” features.
@@
6 Data and code availability
Together with this manuscript, we release the sae-auto-interp library and accompanying scripts to guide the reproduction of these results.
In collaboration with Neuronpedia we also release the interpretations for the Gemma 2 9b 16k residual stream SAEs, the 16k MLP SAEs, the 131k residual stream SAEs and the Llama 3.1 8b 262k residual stream and MLP SAEs. interpretations can be directly downloaded here.
7 Acknowledgements
We would like to acknowledge Lucia Quirke and David Johnston for the work reviewing and discussing the manuscript. We would like to acknowledge Jacob Drori’s role in the initial development of new scoring techniques. We would also like to thank Joseph Bloom, Sam Marks, Can Rager, Jannik Brinkmann and Sam Marks, Neel Nanda, Sarah Schwettmann and Jacob Steinhardt for their suggestions on an earlier version of the work. We are thankful to Open Philanthropy for funding this work. We are grateful to CoreWeave for providing the compute resources.
Appendix A Appendix
A.1 SAE interpretations of a random sentence
In this section, we show some non cherry-picked examples of latent interpretations on a sentence.
Figure A1: SAE latents interpretations for a random sentence. To visualize the latent interpretations produced, we select a sentence taken from the RPJv2 dataset. We selected 4 tokens in different positions in that sentence and filter for latents that are active in different layers. Then we randomly select active latents and their corresponding interpretations to display. We display the detection and fuzzing scores of each interpretation, which indicate how well it explains other examples in the dataset (see Section 3 for details on these scores). The latents selected had high activation, but were not cherry-picked based on interpretations or scores.
A.2 Explainer Prompt
The system prompt if not using Chain of Thought (COT): {mdframed}

### v3

@@ 3.3.1 Detection
One simple approach for scoring interpretations is to ask a language model to identify whether a whole sequence activates a SAE feature given an interpretation. Detection requires few output tokens for each example used in scoring, meaning examples from a wider distribution of feature activation strengths can be used at approximately the same expense. By including non-activating contexts, this method measures both the precision and recall of the interpretation.
Detection is more “forgiving” than simulation insofar as the scorer does not need to localize the feature to a particular token. This means that detection evaluates whether an interpretation correctly identifies the types of contexts that a feature is active on, even if the interpretation does not correctly identify the actual token.
Both this method and fuzzing (described next) can leverage token probabilities to estimate how certain the scorer model is of their classification, and this is an effect that we believe can be to improve the scoring methods. Details on the prompt in Appendix D.1.
@@ 3.3.2 Fuzzing
When interpretations focus on which tokens the feature activates, but not on which contexts, they can still score high on fuzzing, but comparatively lower on detection, specially if they are common tokens. Thus, evaluating an interpretation on both detection and fuzzing can identify whether a model is classifying examples for the correct reason. Details on the prompt on Appendix D.2.
@@ 3.3.3 Surprisal
[…]
The overall surprisal score of the interpretation is given by the AUROC of its information value when viewed as a classifier distinguishing activating from non-activating contexts. Details on the prompt and how to compute the score in Appendix D.3.
@@ 3.3.4 Embedding
Classifying between active and non-active contexts given a certain interpretation can also be seen as using interpretations of features as “queries” that should be able to retrieve relevant “documents”, contexts where the feature is active, between non-relevant “documents”, non-activating contexts. This way, we take a selection of activating and non-activating contexts that embedded by an encoding transformer, and the similarity between the query and the documents is used as a classifier to distinguish between activating and non-activating contexts, and the score is the AUROC.
If the encoding model is small enough - we used a 400M parameter model - this technique is the fastest and opens up the possibility to evaluate a larger fraction of the activation distribution. We have seen that using a larger embedding model - 7B parameter model - didn’t significantly improve the scores, see fig A3, although we believe that this approach was under-investigated. Details on the prompt, on the embedding model and on the way to compute the score in Appendix D.4.
@@ 3.3.5 Intervention scoring
Unlike the above four context-based scores, intervention scoring interprets a feature’s counterfactual impact on model output. We quantify the interpretability of an intervention $I$ with interpretation $\mathbf{z}$ on a distribution of prompts $\pi$ as the average decrease in the scorer’s surprisal about the interpretation when conditioned on text generated with the intervention.
$S=\mathbb{E}_{\mathbf{x}}\Big{[}\mathbb{E}_{\mathbf{i}\sim\mathcal{G}_{I}(\mathbf{x})}[\log p_{\mathcal{M}}(\mathbf{z}|\mathbf{i})]-\mathbb{E}_{\mathbf{g}\sim\mathcal{G}(\mathbf{x})}[\log p_{\mathcal{M}}(\mathbf{z}|\mathbf{g})]\Big{]}$
@@
$\sigma(I;\pi)=\mathbb{E}_{\mathbf{x}\sim\pi}\big{[}D_{KL}(p_{\text{subject}}(\cdot|\mathbf{x})\ ||\ p_{\text{subject},I}(\cdot|\mathbf{x}))\big{]}$
See Appendix G for details on the interpretation and scoring pipeline we use in our intervention experiments.
@@ 4.1 Comparing scoring methods
Input tokens Output tokens
Cost (\$/100k features)
Llama 70b
Cost (\$/100k features)
Claude Sonnet 3.5
@@
When wanting to evaluate the explanation over a larger number of examples, simpler methods than simulation scoring have to be used (Templeton et al., 2024). Simulation scoring is traditionally done only in activating examples, ignoring whether the proposed explanation correctly handles non-activating examples. Simulation scoring is normally done in a small amount of examples per feature due to its high cost, and this is a significant disadvantage that prompted us to investigate more efficient scoring methods.
In Table 1 we compare the amount of tokens used to score a single feature using 100 different contexts, for both fuzzing, detection and two different simulation methods. We can see that fuzzing and detection are at least 5x cheaper than simulation when using the all-at-once (AAO) method described in (Bills et al., 2023), although this method requires access to the log-probabilites of the prompt tokens, which are not accessible from providers of state-of-the-art closed-source models. A researcher wanting to generate simulation scores on the fly, without access to a local model, would have to use the token-by-token (TBT) method, which is 30x more than expensive than both fuzzing and detection. For the same set of intepretations, fuzzing and detection scores given by Llama 70b and Claude Sonnet 3.5 have similar distributions. On the other hand, Claude Sonnet 3.5 produces to higher simulation scores on average than Llama 70b, see Table A9 meaning that we should probably compare the price of fuzzing and detection on Llama 70b to the price of simulation using Claude Sonnet 3.5.
Embedding scoring is even cheaper, requiring 4000 input tokens for 100 different contexts, which at an average of \$0.13 per million tokens would only cost \$50/100k features.
@@
We propose that feature interpretations should be evaluated using more than a single technique when possible, because as discussed in Section 3.3, each of the proposed methods have different failure modes, and a more rich scoring framework can address some of these flaws. Specifically, we find that when computing the correlation of scores computed on the same 800 explanations, we find that while there is a clear positive correlation, some scores disagree – see Appendix D.6 and Tables A1 and A2.
@@
We find fuzzing to be the most correlated with simulation, as fuzzing is mostly measuring how well the explanation can help the model predict which tokens have non-zero activation. Due to the sparsity of SAE activations, simulation scoring is doing a very similar test. On the contrary, we find that simulation has a significantly lower correlation with detection, embedding and surprisal scoring, as they more accurately measure whether a context activates a given feature, instead of which token is active in a given context. When comparing the simulation scores provided by Claude 3.5 Sonnet, all correlations have a slight increase, see Appendix D.6.
@@
We also measure the correlation between our scoring methods and scores given by humans. Humans were tasked to score how an explanation related to a given context, similar to that done in (Templeton et al., 2024). The human evaluators could choose between 4 options, “1 - Explanation not related to the text”, “2 - Explanation vaguely related to the text”, “3 - Incomplete explanation but present in the text”, “4 - Explanation describes the activation of the tokens”. We measure the average score of a feature in at least 5 contexts. In total 700 contexts and 81 feature interpretation. We find that fuzzing as the best Spearman correlation with human scores (0.69), followed by simulation (0.60), and detection (0.59). Surprisal and embedding have 0.34 and 0.32 Spearman correlation respectively.
@@
Due to their relative cheapness compared with simulation scoring, we propose that fuzzing and detection scoring can be used to estimate whether explanations correctly identify on which tokens a feature is most likely to be active, and in which types of contexts this happens. Cheaper methods like embedding can be used to quickly iterate on explanations or to broadly separate features that have bad explanations from those with good explanations, so that they can be refined.
@@ 4.1.1 Intervention scoring
The previously discussed scoring methods are correlational in the sense that they measure whether an interpretation can be used to predict the activation values of a given feature, or distinguish between activating and non-activating examples. Intervention scoring proposes to measure how well a given interpretation can predict the effect of interventions on the corresponding feature. Here we compare correlational interpretations generated with our pipeline and scored with fuzzing, to a set of interventional interpretations scored with intervention scoring. Our hypothesis is that some features will have low correlational scores because their behavior is better explained by their downstream effects than by the contexts where they are active.
In Figure 3, we see that there is a slight negative correlation between the fuzzing score and the intervention score, showing that on average, features with low fuzzing score can be better explained using their downstream effects than features with high fuzzing score. We also find that the distribution of intervention scores of interpretation on a trained SAE is significantly different from the distribution of an interpretation of a randomly initialized SAE or a real SAE with randomly assigned interpretations, supporting the validity of this scoring method.
Figure 3: Intervention scores. Here we present intervention scores (Sec 3.3.5) for SAE features in Gemma 2 9B at layer 32. Left: SAE features are more interpretable than random features, especially when intervening more strongly. Our explainer also produces interpretations that are scored higher than random interpretations. Right: Many features that would normally be uninterpreted when using context-based automatic interpretability are interpretable in terms of their effects on output.
@@ 4.2 Comparing interpretation methods
We use a set of 500+ features as a testbed to measure the effects of design choices and hyperparameters on interpretation quality. Each feature is scored using 100 activating and 100 non-activating examples. The activating examples are chosen via stratified sampling such that there are always 10 examples from each of the 10 deciles of the activation distribution. We evaluate interpretation quality using fuzzing, detection and embedding scores, as those were both quick to compute and easy to interpret. We expect this mix of scores reflect the extent to which the proposed interpretation is valid over both activating and non-activating examples.
@@
We find that the interpretations generated by considering only the activating contexts are significantly different from those generated by selecting activating contexts from the whole distribution. Interpretations based on top examples have higher specificity, but lower sensitivity. In fact, we observe that the sensitivity depends on how the degree of activation of each example shown, and that this sensitivity decreases faster for interpretation based on top examples. Interpretations generated from top activating examples are more concise and specific interpretations but fail to capture the whole distribution, while explanations based on examples sampled from the whole distribution lead to interpretations that are too broad, see Appendix C. This effect would not be seen if the scoring were done on just the most activating examples, underscoring a problem with current auto-interpretability evaluations, which produce interpretations using top activating examples and evaluate them on a small subset of the activation distribution.
@@
Increasing the size of the explainer model increases the scores of the interpretations, but we don’t find the explanations generated by Claude Sonnet 3.5 to have much higher scores than those generated by Llama 3.1 70b, see A9, and both are similar to those generated by a human. Not surprisingly, we also see that even for simpler scores like fuzzing and detection, using a smaller scorer model leads to lower scores. It is possible the explanations generated by Claude would be better than those generated by Llama 3.1 70b had we optimized the prompting techniques for that model.
@@
Showing the explainer model a larger number of examples leads to slightly higher scores (Table A7). We find it doesn’t matter much whether we use the same dataset used to train the SAEs (Table A3) or whether we slightly change the size of the contexts shown (Table A5). Using COT, at least when generating interpretations using Llama 70b, does not significantly increase their quality (Table A4) while significantly increasing the compute and time required to generate them. For this reason, we have not used it for our main experiments.
@@
We find that SAEs with more features have higher scores, and scores that are significantly higher than those of neurons. Neurons are more interpretable if made sparser by only considering the top $k$ most activated neurons on a given token, but still significantly underperform SAEs in our tests (Table A10). The location of the SAE matters; residual stream SAEs have slightly better scores than ones trained on MLP outputs (Table A10). We also observe that earlier layers have lower overall scores, but that the scores across model depth remain constant after the first few layers (Fig A5).
@@ 5 Conclusion
Explaining the features of SAEs trained on cutting-edge LLMs is a computationally demanding task, requiring scalable methods to both generate interpretations and assess their quality. We addressed issues with the conventional simulation-based scoring and introduced five new scoring techniques, each with distinct strengths and limitations. These methods allowed us to explore the “prompt design space” for generating effective interpretations and propose practical guidelines.
Additionally, although current scoring methods do not account for interpretation length, we believe shorter interpretations are generally more useful and will incorporate this in future metrics. Some scoring methods also require further refinement, particularly in selecting non-activating examples to improve evaluation. We also investigated a method to generate and score interpretations of features based on their downstream effect and shown that low score “correlational” scores could be due to the existence of “output” features.
@@
Data and Code Availability
Together with this manuscript, we release the delphi library and accompanying scripts to guide the reproduction of these results.
In collaboration with Neuronpedia we also release the explanations for the Gemma 2 9b 16k residual stream SAEs, the 16k MLP SAEs, the 131k residual stream SAEs and the Llama 3.1 8b 262k residual stream and MLP SAEs. Explanations can be directly downloaded here.
Acknowledgements
We would like to acknowledge Lucia Quirke and David Johnston for the work reviewing and discussing the manuscript. We would like to acknowledge Jacob Drori’s role in the initial development of new scoring techniques. We would also like to thank Joseph Bloom, Sam Marks, Can Rager, Jannik Brinkmann and Sam Marks, Neel Nanda, Sarah Schwettmann and Jacob Steinhardt for their discussions and suggestions in an earlier version of the work. We are thankful to Open Philanthropy for funding this work. We are grateful to CoreWeave for providing the compute resources.
Impact Statement
This paper presents work whose goal is to advance the field of Mechanistic Interpretability. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A SAE interpretations of a random sentence
In this section, we show some non cherry-picked examples of feature interpretations on a sentence.
Figure A1: SAE features interpretations for a random sentence. To visualize the feature interpretations produced, we select a sentence taken from the RPJv2 dataset. We selected 4 tokens in different positions in that sentence and filter for features that are active in different layers. Then we randomly select active features and their corresponding interpretations to display. We display the detection and fuzzing scores of each interpretation, which indicate how well it explains other examples in the dataset (see Section 3 for details on these scores). The features selected had high activation, but were not cherry-picked based on interpretations or scores.
Appendix B Explainer Prompt
The system prompt if not using Chain of Thought (COT):